OpenStudy (anonymous):

help? medals given given the function f (x) = x^3 + 2x^2 - 3x - 5, what is the resulting function when f (x) is shifted to the right one unit? pleasee.

OpenStudy (anonymous):

someone!!

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Directrix @Loser66

OpenStudy (loser66):

so fun, I've just received your 1-hour-sending message. I am sorry. The net is so bad. To your problem, just replace x = x-1 and simplify it so , f(x-1) = (x-1)^3 +2(x-1)^2-3(x-1) - 5 = x^3 -3x^2+3x -1 +2x^2 -4x +2 -3x +3 -5 = x^3 -x^2 -4x -1